OpenStudy (anonymous):

tan^2x-1/1-tanx I need to get it down to one function with no fractions. I got to (sin^2x/cosx)-cosx but I don't know where to go from there or if that is right.

OpenStudy (anonymous):

Use difference of squares.

OpenStudy (anonymous):

\[ a^2-b^2 = (a-b)(a+b) \]

OpenStudy (anonymous):

could you show me because I don't see how that would work in this problem

OpenStudy (anonymous):

\[ a=\tan(x),b=1 \]

OpenStudy (anonymous):

\[ \tan^2(x)-1 = (\tan(x)-1)(\tan(x)+1) \]

OpenStudy (anonymous):

\[ 1-\tan(x) = -(\tan(x)-1) \]

OpenStudy (anonymous):

idk if this will help but 1+tan x=cosx

OpenStudy (anonymous):

....

OpenStudy (anonymous):

\[ \frac{\tan^2x-1}{1-\tan x} = \frac{(\tan x-1)(\tan x+1)}{-(\tan x-1)} \]

OpenStudy (anonymous):

I've tried to get help with this problem before but I put part of the question wrong. I think the other guy was on the right track but was wrong cause I squared something that shouldn't have been squared

OpenStudy (anonymous):

ooooo that should work